Find the value of h if the slopes of the lines represented by the 6x^2 + 2hxy + y^2 are in the ratio 1:2

Grade:12

1 Answers

saiomkar kandukuri
11 years ago

equation for pair of st lines passing through origin:

(y-m1x)*(y-m2x)=0   where m1 ,m2 are slopes of two lines

question is   6X^2 +2hxy +y^2=0

m1  /m2 = 1/2

asper equation of pair of st lines passing through origin

m1*m2 =6 ... m1/m2 =1/2.... -(m1 +m2)=2h

solve the above three equations and get the answer.......